OpenStudy (anonymous):

Angle DEF and angle FEG are complementary. The measure of angle DEF = (3x-4)º and the measure of angle FEG = (5x+6)º. What is the measure of angle DEF?

OpenStudy (anonymous):

I keep getting 26 :/

OpenStudy (anonymous):

do you know complementary angles and what do their sum equals to ?

OpenStudy (anonymous):

Yeahh complementary = 90 degrees

OpenStudy (anonymous):

correct !! 3x-4+5x+6=90 now solve for x

OpenStudy (anonymous):

Thats what I was doing.. 3x nd 5x are like terms. So you combine them right? nd get 8x

OpenStudy (anonymous):

yes 8x+2=90

OpenStudy (anonymous):

what you got for x

OpenStudy (anonymous):

Thanks! I was combining 4 and 6.. nd getting 8x + 10 = 90. But i see now (:

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Whts the answer